# Short-horizon Duesenberry Equilibrium

### 17 Mar 2026 v1

@@ A.1. Semimartingale property of aggregated processes
Let $f\colon[0,\infty)\times\mathbb{D}\to\mathbb{R}$ be a continuous function with continuous partial derivatives $\partial^{2}f/\partial x_{i}\partial x_{j}$ and $\partial f/\partial t$, for $i,j=1\cdots d$ ($f\in C^{1,2}([0,\infty)\times\mathbb{D}$), where $\mathbb{D}\subset\,\mathbb{R}^{d}$ is an open subset. Suppose that $\varphi$ is consistent semimartingale process with values in $\mathbb{D}$ with stochastic representation (2) where $\rho$ and $\varrho$ are functions of class $C^{m,\chi}$ where $m=0$ and $\chi=1$ or $m\geq 1$, and $\chi>0$. Assume that $\mu$ is a probability distribution on $\mathbb{D}$ with a continuous density $g(x)$. Assume that for any $s<t<\infty$:
@@ 2. $i$
    $\partial f/\partial t$, $\sum_{i}\rho^{i}\partial f/\partial x_{i}$, $\sum_{i}\varrho^{i,k}\partial f/\partial x^{i}$, $\sum_{i,j}\varrho^{i,j}\partial^{2}f/\partial x_{i}\partial x_{j}$ are Lipschitz continuous for all $k$, where $\varrho^{k}=(\varrho^{1,k},\cdots,\varrho^{d,k})$ and $\rho=(\rho^{1},\cdots,\rho^{d})$ for $k=1,\cdots,n$.
@@
    $E[\int_{\mathbb{D}}|\varphi^{i}_{s,t}(x)|d\mu(x)]<\infty$ for $i=1,\cdots,d$.
@@
$\int_{\mathbb{D}}f(t,\varphi_{s,t}(x))\,d\mu(x)=\\ \int_{\mathbb{D}}f(s,x)g(x)\,dx+\int_{s}^{t}\int_{\mathbb{D}}\frac{\partial}{\partial t}f(u,\varphi_{s,u}(x))g(x)\,dx\,du+\\ \int_{s}^{t}\sum_{i}\int_{\mathbb{D}}\frac{\partial}{\partial x_{i}}f(u,\varphi_{s,u}(x))\rho^{i}(u,\varphi_{s,u}(x))g(x)dx\,du+\\ \frac{1}{2}\int_{s}^{t}\sum_{i,j}\int_{\mathbb{D}}\frac{\partial^{2}}{\partial x_{i}\partial x_{j}}f\left(u,\varphi_{s,u}(x))\varrho^{i,j}(u,\varphi_{s,u}(x)\right)g(x)\,dx\,du+\\ \sum_{k}\int_{s}^{t}\sum_{i}\int_{\mathbb{D}}\frac{\partial f}{\partial x_{i}}(u,\varphi_{s,u}(x))\varrho^{i,k}(u,\varphi_{s,u}(x))g(x)\,dx\,dW_{s}^{k}(u).$ (61)
@@
$\int_{\mathbb{D}}f(t,\varphi_{s,t}(x))\,d\mu(x)=\\ \int_{\bar{U}}f(t,\varphi_{s,t}(x))g(x)\,dm=\lim_{|P_{r}|\to 0}\sum_{x\in P_{r}}f(t,\varphi_{s,t}(x^{\star}))g(x^{\star})m([x,x^{+}))=\\ \lim_{|P_{r}|\to 0}\sum_{x\in P_{r}}f(t,x^{\star})g(x^{\star})m([x,x^{+}))+\\ \lim_{|P_{r}|\to 0}\int_{s}^{t}\sum_{x\in P_{r}}\frac{\partial}{\partial t}f(u,\varphi_{s,u}(x^{\star}))g(x^{\star})m([x,x^{+}))\,du+\\ \lim_{|P_{r}|\to 0}\int_{s}^{t}\sum_{i}\sum_{x\in P_{r}}\frac{\partial}{\partial x_{i}}f(u,\varphi_{s,u}(x^{\star}))\rho^{i}(u,\varphi_{s,u}(x^{\star}))g(x^{\star})m([x,x^{+}))\,du+\\ \lim_{|P_{r}|\to 0}\frac{1}{2}\int_{s}^{t}\sum_{i,j}\sum_{x\in P_{r}}\frac{\partial^{2}}{\partial x_{i}\partial x_{j}}f\left(u,\varphi_{s,u}(x^{\star}))\varrho^{i,j}(u,\varphi_{s,u}(x^{\star})\right)g(x^{\star})m([x,x^{+}))\,du+\\ \lim_{|P_{r}|\to 0}\sum_{k}\int_{s}^{t}\sum_{i}\sum_{x\in P_{r}}\frac{\partial f}{\partial x_{i}}(u,\varphi_{s,u}(x^{\star}))\varrho^{i,k}(u,\varphi_{s,u}(x^{\star}))g(x^{\star})m([x,x^{+}))\,dW_{s}^{k}(u),$ (62)

### 18 Mar 2026 v2

@@ A.1. Semimartingale property of aggregated processes
Let $f\colon[0,\infty)\times\mathbb{D}\to\mathbb{R}$ be a continuous function with continuous partial derivatives $\partial^{2}f/\partial x_{i}\partial x_{j}$ and $\partial f/\partial t$, for $i,j=1,\dots,d$ (that is, $f\in C^{1,2}([0,\infty)\times\mathbb{D})$), where $\mathbb{D}\subset\mathbb{R}^{d}$ is an open subset. Suppose that $\varphi$ is a consistent semimartingale process with values in $\mathbb{D}$ and stochastic representation (2), where $\rho$ and $\varrho$ are functions of class $C^{m,\chi}$ with $m=0$ and $\chi=1$ or $m\geq 1$, and $\chi>0$. Assume that $\mu$ is a probability distribution on $\mathbb{D}$ with continuous density $g(x)$. Define
$a^{ij}(t,x)\coloneqq\sum_{k=1}^{n}\varrho^{i,k}(t,x)\varrho^{j,k}(t,x),\qquad i,j=1,\dots,d.$
Assume that for any $s<t<\infty$:
@@ 2. $i$
    $\partial f/\partial t$, $\sum_{i}\rho^{i}\,\partial f/\partial x_{i}$, $\sum_{i}\varrho^{i,k}\,\partial f/\partial x_{i}$ for every $k=1,\dots,n$, and $\sum_{i,j}a^{ij}\,\partial^{2}f/\partial x_{i}\partial x_{j}$ are Lipschitz continuous.
@@
    $E\!\left[\int_{\mathbb{D}}|\varphi^{i}_{s,t}(x)|\,d\mu(x)\right]<\infty$ for $i=1,\dots,d$.
@@
$\int_{\mathbb{D}}f(t,\varphi_{s,t}(x))\,d\mu(x)=\int_{\mathbb{D}}f(s,x)g(x)\,dx\\ +\int_{s}^{t}\int_{\mathbb{D}}\frac{\partial}{\partial t}f(u,\varphi_{s,u}(x))g(x)\,dx\,du\\ +\int_{s}^{t}\sum_{i}\int_{\mathbb{D}}\frac{\partial}{\partial x_{i}}f(u,\varphi_{s,u}(x))\rho^{i}(\varphi_{s,u}(x))g(x)\,dx\,du\\ +\frac{1}{2}\int_{s}^{t}\sum_{i,j}\int_{\mathbb{D}}\frac{\partial^{2}}{\partial x_{i}\partial x_{j}}f(u,\varphi_{s,u}(x))a^{ij}(\varphi_{s,u}(x))g(x)\,dx\,du\\ +\sum_{k}\int_{s}^{t}\sum_{i}\int_{\mathbb{D}}\frac{\partial f}{\partial x_{i}}(u,\varphi_{s,u}(x))\varrho^{i,k}(\varphi_{s,u}(x))g(x)\,dx\,dW_{s}^{k}(u).$ (61)
@@
$\int_{\mathbb{D}}f(t,\varphi_{s,t}(x))\,d\mu(x)=\\ \int_{\bar{U}}f(t,\varphi_{s,t}(x))g(x)\,dm=\lim_{|P_{r}|\to 0}\sum_{x\in P_{r}}f(t,\varphi_{s,t}(x^{\star}))g(x^{\star})m([x,x^{+}))=\\ \lim_{|P_{r}|\to 0}\sum_{x\in P_{r}}f(t,x^{\star})g(x^{\star})m([x,x^{+}))+\\ \lim_{|P_{r}|\to 0}\int_{s}^{t}\sum_{x\in P_{r}}\frac{\partial}{\partial t}f(u,\varphi_{s,u}(x^{\star}))g(x^{\star})m([x,x^{+}))\,du+\\ \lim_{|P_{r}|\to 0}\int_{s}^{t}\sum_{i}\sum_{x\in P_{r}}\frac{\partial}{\partial x_{i}}f(u,\varphi_{s,u}(x^{\star}))\rho^{i}(\varphi_{s,u}(x^{\star}))g(x^{\star})m([x,x^{+}))\,du+\\ \lim_{|P_{r}|\to 0}\frac{1}{2}\int_{s}^{t}\sum_{i,j}\sum_{x\in P_{r}}\frac{\partial^{2}}{\partial x_{i}\partial x_{j}}f\bigl(u,\varphi_{s,u}(x^{\star})\bigr)a^{ij}\bigl(\varphi_{s,u}(x^{\star})\bigr)g(x^{\star})m([x,x^{+}))\,du+\\ \lim_{|P_{r}|\to 0}\sum_{k}\int_{s}^{t}\sum_{i}\sum_{x\in P_{r}}\frac{\partial f}{\partial x_{i}}(u,\varphi_{s,u}(x^{\star}))\varrho^{i,k}(\varphi_{s,u}(x^{\star}))g(x^{\star})m([x,x^{+}))\,dW_{s}^{k}(u).$ (62)